OpenStudy (anonymous):

3 sin theta = sin theta -1

OpenStudy (anonymous):

@teddybear104

OpenStudy (anonymous):

wats up

OpenStudy (anonymous):

hey I need help badly

OpenStudy (anonymous):

go for it

OpenStudy (anonymous):

I don't know how to simplify the equation above

OpenStudy (anonymous):

I will give you a medal if you can help

OpenStudy (anonymous):

@domebotnos

OpenStudy (domebotnos):

sorry :^( i don't know this yet

OpenStudy (anonymous):

please help me I need to understand this and I don't

OpenStudy (tkhunny):

Unclear what is wanted. \(3\sin(\theta) = \sin(\theta) - 1\) Is that it? Can you solve... 3x = x-1?

OpenStudy (anonymous):

yes and yes

OpenStudy (tkhunny):

Okay, what is the solution to the little x-problem?

OpenStudy (anonymous):

x=-1/2

OpenStudy (tkhunny):

So, why wouldn't the solution to the original problem be \(\sin(\theta) = -1/2\)?

OpenStudy (anonymous):

here are the options